# Taming systematics in distance and inclination measurements with gravitational waves: role of the detector network and higher-order modes

### v1

@@ IV.3.5 Next-to-unity mass ratio binaries II: the effects of biased estimation of the mass ratio
To investigate this aspect, we repeated the same face-on $36\,M_{\odot}$$29\,M_{\odot}$ injection at $\hat{\rho}=100$, but using a narrower prior on the mass ratio. In Fig. 14 we compare the 90% $d_{L}$$\iota$ credible regions for the two configurations, with both injection and recovery performed with IMRPhenomXHM. In Tab. 8 we instead summarize the main results, which witness slight improvements when narrowing the prior on mass ratio. We add that the bias worsens as the SNR increases: for $\hat{\rho}=20$, the parameter estimation provides a mass ratio $q=0.83^{+0.15}_{-0.21}$. The increased SNR thus eliminates mass ratio samples closer to the injected value.
@@ V Discussion and conclusions
Figure 15: Left panel: the overlap between a fixed linearly polarized waveform with mass ratio $q\sim 0.8$, indicated by the red vertical line, and waveforms with the same parameters but with varying mass ratio, according to Eq. (33). The chirp mass is fixed at $\mathcal{M}=28.1\,M_{\odot}$. Right panel: similar to the left panle, but the fixed mass ratio is $q=0.1$, while the chirp mass is fixed at $\mathcal{M}=14.8\,M_{\odot}$ and the other parameters follow the left table of Tab. 1. All waveforms are generated with IMRPhenomXHM.
@@
Building upon this detailed study of higher-order modes, future research will investigate the impact of precession [71]. This analysis presents its own set of challenges, primarily due to the longer timescale of precession relative to orbital variations [87], requiring observations spanning many orbital cycles. 3G detectors will be crucial in this endeavor, enabling the observation of signals from lower frequencies and providing access to longer data stretches. Similar to our findings with higher-order modes, the effects of precession are expected to be more prominent in nearly edge-on binaries [88] and low-mass-ratio systems. Optimal combinations of SNR, spin, and mass ratio are anticipated to yield sufficient accuracy and precision in parameter estimation.
Finally, we emphasize that this work mostly assumed ET in its standard triangular configuration, finding preliminary evidence that the degeneracy will be reduced by a 2L design, mostly due to the larger SNR. Future investigations should explore how alternative ET designs [64, 72] might influence the systematic effects detailed herein. , Regardless of the specific ET design, the results will largely depend on the nature of the source. While BBHs with high-order modes can, in principle, help break the degeneracy, omitting these modes will reintroduce it. This will make the case of BNSs without an electromagnetic counterpart particularly challenging, as high-order modes and precession are generally negligible.

### v2

@@ IV.3.5 Next-to-unity mass ratio binaries II: the effects of biased estimation of the mass ratio
To investigate this aspect, we repeated the same face-on $36\,M_{\odot}$$29\,M_{\odot}$ injection at $\hat{\rho}=100$, but using a narrower prior on the mass ratio. In Fig. 14 we compare the 90% $d_{L}$$\iota$ credible regions for the two configurations, with both injection and recovery performed with IMRPhenomXHM. In Table 8 we instead summarize the main results, which witness slight improvements when narrowing the prior on mass ratio. We add that the bias worsens as the SNR increases: for $\hat{\rho}=20$, the parameter estimation provides a mass ratio $q=0.83^{+0.15}_{-0.21}$. The increased SNR thus eliminates mass ratio samples closer to the injected value.
@@ V Discussion and conclusions
Figure 15: Left panel: the overlap between a fixed linearly polarized waveform with mass ratio $q\sim 0.8$, indicated by the red vertical line, and waveforms with the same parameters but with varying mass ratio, according to Eq. (33). The chirp mass is fixed at $\mathcal{M}=28.1\,M_{\odot}$. Right panel: similar to the left panle, but the fixed mass ratio is $q=0.1$, while the chirp mass is fixed at $\mathcal{M}=14.8\,M_{\odot}$ and the other parameters follow the left table of Table 1. All waveforms are generated with IMRPhenomXHM.
@@
Building upon this detailed study of higher-order modes, future research will investigate the impact of precession [71]. This analysis presents its own set of challenges, primarily due to the longer timescale of precession relative to orbital variations [87], requiring observations spanning many orbital cycles. 3G detectors will be crucial in this endeavor, enabling the observation of signals from lower frequencies and providing access to longer data stretches. Similar to our findings with higher-order modes, the effects of precession are expected to be more prominent in nearly edge-on binaries [88] and low-mass-ratio systems. Optimal combinations of SNR, spin, and mass ratio are anticipated to yield sufficient effectiveness in the parameter estimation.
Finally, we emphasize that this work mostly assumed ET in its standard triangular configuration, finding preliminary evidence that the degeneracy will be reduced by a 2L design, mostly due to the larger SNR. Future investigations should explore how alternative ET designs [64, 72] might influence the systematic effects detailed herein. Regardless of the specific ET design, the results will largely depend on the nature of the source. While BBHs with high-order modes can, in principle, help break the degeneracy, omitting these modes will reintroduce it. This will make the case of BNSs without an electromagnetic counterpart particularly challenging, as higher-order modes and precession are generally negligible.